How many grams of HgO must react to produce 3.75 moles of O2?
1
Expert's answer
2013-02-15T05:43:31-0500
2HgO = 2Hg + O2

According to the chemical equation the amount of HgO is double of O2 produced.

n(HgO) = 2n(O2) = 2 x 3.75 = 7.50 moles

Molar weight of HgO = 200.59 + 16.00 = 216.59 (g/mol)
Molar weights of atoms Hg and O were taken from the periodic table of elements.

The mass of HgO needed = amount of HgO (moles) x molar weight (g/mol)

m = 7.50 moles x 216.59 (g/moles) = 1624.42 g

Answer:

m(HgO) = 1624.42 g